OpenStudy (anonymous):

find the 9th term in the following geometric sequence.. round your answer to the nearest tenths.. 5,8.5,14.45,24.565....

OpenStudy (anonymous):

common ratioo = 8.5 / 5 = 1.7 9th term = ar^ 9-1 = 5 * 1.7 ^ 8 sorry i havent got my calculator to work this out

OpenStudy (anonymous):

348.8